OpenStudy (anonymous):

ill give a medal if someone helps with this question 7th grade math

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

! ♥

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

Dakota divided the rational numbers as follows: two-fifths divided by two and three-tenths equals what value? Step one: equals two-fifths divided by twenty-three over ten Step two: equals five-halves times twenty-three over ten Step three: equals one-half times twenty-three over two Step four: twenty-three fourths equals five and three-fourths In which step was Dakota's error? Step 1 Step 2 Step 3 Step 4

OpenStudy (anonymous):

Give me a sec, I'll help

OpenStudy (anonymous):

Step 4 is wrong because she multiplied the denominator, you don't multiply the denominator when multiplying fractions.

OpenStudy (anonymous):

so the correct answer would be 11 and 1/2, but you don't need that, step 4 is the correct answer for your question

OpenStudy (anonymous):

can u help wth 3 more

OpenStudy (anonymous):

I was just about to ask if you had anymore lol

OpenStudy (anonymous):

Yes, I will

OpenStudy (anonymous):

its a really easy one

OpenStudy (anonymous):

Find the product: three-fourths times twelve. one-sixteenth one-ninth 6 9

OpenStudy (anonymous):

9

OpenStudy (anonymous):

12/4 is 3 so 3/4 of 12 is 9

OpenStudy (anonymous):

four and two thirds divided by one and one-sixth equals what value? 2 4 four and one-ninth five and four-ninths

OpenStudy (anonymous):

Start by making them both improper fractions

OpenStudy (anonymous):

14/3 and 7/6

OpenStudy (anonymous):

question 1 2/5 divided by 2 3/10 = 2/5 divided by 23/10 = 2/5 * 10/23 step 2 is wrong......you " flip " what you are dividing by

OpenStudy (anonymous):

So then you multiply by the reciprocal

OpenStudy (anonymous):

@kelliegirl33 Your right, I forgot that you do multiply the denominator of a fraction. So @i_need_help123 your answer is step 2

OpenStudy (anonymous):

yes....you flip what you are dividing by (2/5) / (23/10) = 2/5 * 10/23 NOT 5/2 * 23/10

OpenStudy (anonymous):

So question 3, 14/3 x 6/7

OpenStudy (anonymous):

Sorry, thank you for the correction @kelliegirl33

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

so you get 84/21 or 4/1 which is 4

OpenStudy (anonymous):

So your answer is 4

OpenStudy (anonymous):

4 2/3 divided by 1 1/6 = (14/3) / (7/6) = 14/3 * 6/7 ...yes you got that right

OpenStudy (anonymous):

You have one more right @i_need_help123?

OpenStudy (anonymous):

Gary has a length of string for making a yo-yo that is fifty-seven and one-half inches long. He wants to cut that in half to get the proper length. Which expression can be used to find the length of each piece of string? one-half times fifty-seven and one-half two times fifty-seven and one-half fifty-seven and one-half divided by one-half two divided by fifty-seven and one-half

OpenStudy (anonymous):

Actually, both a and c are correct

OpenStudy (anonymous):

well there has 2 be one

OpenStudy (anonymous):

I think they would want you to go with c, because that makes more sense with the question

OpenStudy (anonymous):

Wait, no, just kidding, the answer is a

OpenStudy (anonymous):

lets see what grade I get

OpenStudy (anonymous):

it's a!

OpenStudy (anonymous):

because it would need to be divided by 2, not 1/2

OpenStudy (anonymous):

but multiplying by 1/2 would work

OpenStudy (anonymous):

its an A woooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo thx so super much ill give you a medal and ill fan you!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):